OpenStudy (anonymous):

What is the r value of the following geometric sequence? -81, 54, -36, 24,…

OpenStudy (anonymous):

\[-81\times r=54\] \[54\times r=-36\] \[-36\times r=24\] and so on. If there is indeed a common ratio \(r\), the solution to each equation should be the same.

OpenStudy (anonymous):

my options are -3/4 -5/6 -5/9 -2/3